OpenStudy (anonymous):

Brandon uses 2over3 of a jar of peanut butter each week. He wants to know how many weeks it will take to use 4 jars. Use the model to find 4 ÷ 2over3, the number of weeks it will take.

OpenStudy (anonymous):

Help plz

OpenStudy (anonymous):

I'm thinking... don't know if I know how to do this or not.... what is this 'model' thing?

OpenStudy (anonymous):

ok take your time

OpenStudy (whpalmer4):

Why don't you try making a graph showing weeks along the x-axis, and jars of peanut butter consumed along the y-axis. Draw a horizontal line at 4 jars, and see where the two lines intersect. Drop a vertical line from that point to the x-axis and read off the number of weeks it takes...

OpenStudy (anonymous):

okay, 4 divided by 2/3 is just 4 times 3/2, so you multiply 4 times 3 and divide by 2, giving you 6

OpenStudy (anonymous):

...ir you could do that...

OpenStudy (whpalmer4):

Yeah, it's an interesting question what is meant by "the model"...

OpenStudy (anonymous):

so the answer is 6?

OpenStudy (whpalmer4):

1 week 2/3 2 week 2/3 + 2/3 = 4/3 3 week 2/3 + 4/3 = 6/3 = 2 another 3 weeks to use up another 2 so 6 weeks to use 4

OpenStudy (anonymous):

yes, its 6

OpenStudy (whpalmer4):

Probably after 4 jars of peanut butter in 6 weeks, you'll be ready to switch to something else :-)

OpenStudy (anonymous):

i will give both of you a medal for helping i will have a other thing for for some one to answer you could if you want.

OpenStudy (anonymous):

wait i can't sorry :(

OpenStudy (anonymous):

awesome, thanks! and yeah, hit me with another, ill see what I can do for you :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i only gave one of you a meadl

OpenStudy (whpalmer4):

you can only award 1 medal per posting

OpenStudy (anonymous):

hears the other one:Tristan walks 3over4 of a mile to school every day. He wants to find how long it will take him to walk 6 miles. Use the model to find 6 ÷ 3over4, the number of days it will take Tristan to walk 6 miles.

OpenStudy (anonymous):

its pretty much the same, taking the inverse of the second figure. its 6 divided by 3/4 which is the same as 6 times 4/3 which simplifies to 8

OpenStudy (anonymous):

oh and on the medal thing, if you really wanted to, you could repost your second question in a new post and whichever one of us you didn't give the first one to could take the second one